Circular No. 7698

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q6 (NEAT) S. H. Pravdo, E. F. Helin, and K. J. Lawrence, Jet Propulsion Labora-

tory, report the discovery of another comet on CCD images taken with the 1.2-m Schmidt telescope at Palomar in the course of the NEAT program; their discovery image on Aug. 28 (given below) shows a central nebulosity of diameter ∼ 3′′ and a tail ∼ 10′′ long toward the west-southwest. Other reported physical descriptions from CCD images include: Aug. 28.98 UT, diffuse (M. Tichý, Kleť, 0.57-m reflector); 29.05, diffuse object (m_{1} = 16.4) with a bright nucleus, m_{2} = 17.9 (L. Šarounová and M. Wolf, Ondřejov, 0.65-m reflector); 29.38, coma diameter ∼ 10′′ (K. Smalley, Louisburg, KS, 0.75-m reflector); 29.44, well-condensed coma, broad tail 15′′ long in p.a. 240^{o} (D. Balam, Dominion Astrophysical Observatory, 1.82-m Plaskett telescope); 29.47, fuzzy with a hint of a bulge to the southwest (P. J. Shelus, McDonald Observatory, 0.76-m reflector).

2001 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 28.41375  3^{h}35^{m}29.^{s}10  +9^{o}32′40′′.2  17.6

Additional astrometry and the following preliminary parabolic orbital elements (from 25 observations, Aug. 28–29) appear on MPEC 2001-Q70:

T = 2001 Nov. 3.898 TT  ω = 38.^{o}984
Ω = 23.312 }2000.0
q = 1.46749 AU  i = 57.518
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 29 3^{h}35.^{m}92 + 9^{o}54.^{′}8 1.230 1.740 101.^{o}5 34.^{o}7 16.4
Sept. 3 3 39.46 +13 17.5 1.141 1.705 104.7 34.9 16.1
8 3 42.46 +17 09.7 1.056 1.671 108.1 35.0 15.8
13 3 44.81 +21 36.7 0.978 1.639 111.4 34.9 15.6
18 3 46.29 +26 43.4 0.906 1.609 114.7 34.6 15.4
23 3 46.61 +32 33.4 0.843 1.582 117.6 34.2 15.1
28 3 45.33 +39 07.0 0.792 1.557 120.1 33.8 14.9
Oct. 3 3 41.72 +46 19.1 0.753 1.535 121.6 33.7 14.7
8 3 34.48 +53 57.0 0.728 1.516 122.0 34.0 14.6

V4739 SAGITTARII Photometry by A. C. Gilmore, Mt. John (IAUC 7695): Aug. 29.349

UT, V = 9.94, U–B = −0.85, B–V = −0.03, V –R = +1.37, V –I = +1.45.

2001 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green